one side of right traingle is 12 cm long. The hyphotenuse is 3 cm more than twice as long as the other side. What are the dimensions of the rectangle?
Answer by mananth(16946)

one side = 12 cm
other side =x
hypotenuse = 2x+3
By Hypotenuse theorem
(2x+3)^2=x^2+12^2
4x^2+12x+9=x^2+144
3x^2+12x-135=0
/3
x^2+4x-45=0
x^2+9x-5x-45=0
x(x+9)-5(x+9)=0
(x+9)(x-5)=0
therefore x= -9 OR 5
ignore negative value
x=5 cm the other side
Hypotenuse = 2x+3.
Plug x=6
we get hypotenuse = 13
The dimensions are , 5,12,13